A bath contains 100kg of water at 60°C. Hot and cold taps are then turned on to deliver 20kg per minute each at temperatures 70°C and 10°C respectively. How long will it be before the temperature in the bath dropped to 45°C?
1
Expert's answer
2020-05-27T10:17:42-0400
"Mc(T_0-T)=\frac{dm}{dt}c(T_H+T_C)t"

"100(100-45)=20(70+10)t"

"t=3.4375\ min"